A conical pendulum has a mass of 1kg and the massless string has a length of 1 meter. What is the period of the pendulum's motion if the string makes an angle of 12 degrees with the y-axis? Thank you in advance!


1
Expert's answer
2021-04-15T10:44:18-0400

The period is


"T=2\pi\sqrt{\frac{l\cos\theta}{g}}=1.99\text{ s}."